OpenStudy (anonymous):

Find the x- and y-intercepts and graph the equation plotting the intercepts. Please show all of your work. Submit your graph to the Dropbox. -2x+5y=-15

OpenStudy (accessdenied):

x-intercept is when y = 0 -2x + 5(0) = -15 -2x = -15 x = 15/2 ~ 7.5 (7.5, 0) y-intercept is when x = 0 -2(0)+5y = -15 5y = -15 y = -3 (0, -3) Graph is something like... 

OpenStudy (anonymous):

How can I copy and paste the graph?

OpenStudy (accessdenied):

http://i42.tinypic.com/2zfib2o.png could probably copy paste that (points are in red, line in black)